# Multi-scale Temporal Prediction via Incremental Generation and Multi-agent Collaboration

### v2

@@ Abstract
Accurate temporal prediction is the bridge between comprehensive scene understanding and embodied artificial intelligence. However, predicting multiple fine-grained states of scene at multiple temporal scales is difficult for vision-language models. We formalize the Multi‐Scale Temporal Prediction (MSTP) task in general and surgical scene by decomposing multi‐scale into two orthogonal dimensions: the temporal scale, forecasting states of human and surgery at varying look‐ahead intervals, and the state scale, modeling a hierarchy of states in general and surgical scene. For instance in general scene, states of contacting relationship are finer-grained than states of spatial relationship. For instance in surgical scene, medium‐level steps are finer‐grained than high‐level phases yet remain constrained by their encompassing phase. To support this unified task, we introduce the first MSTP Benchmark, featuring synchronized annotations across multiple state scales and temporal scales. We further propose a novel method, Incremental Generation and Multi‐agent Collaboration (IG-MC), which integrates two key innovations. Firstly, we propose an plug-and-play incremental generation to keep high-quality temporal prediction that continuously synthesizes up-to-date visual previews at expanding temporal scales to inform multiple decision-making agents, ensuring decision content and generated visuals remain synchronized and preventing performance degradation as look‐ahead intervals lengthen. Secondly, we propose a decision‐driven multi‐agent collaboration framework for multiple states prediction, comprising generation, initiation, and multi‐state assessment agents that dynamically triggers and evaluates prediction cycles to balance global coherence and local fidelity. Extensive experiments on the MSTP Benchmark in general and surgical scene show that IG‐MC is a generalizable plug-and-play method for MSTP, demonstrating the effectiveness of incremental generation and the stability of decision‐driven multi‐agent collaboration.
@@ 1 Introduction
In a parallel vein, recent advances in artificial intelligence [19, 20, 21, 22] have significantly improved analysis and prediction of healthcare robot, achieving human-level performance in tasks such as phase segmentation [23] and instrument-tissue interaction detection [24] through convolutional [25, 26, 27], recurrent [28], and transformer-based [29, 30, 31] architectures, as well as temporal [23, 32, 33] and diffusion models [34, 35]. Undoubtedly, these approaches made an initial breakthrough, enabling faithful and effective surgical problem-solving. However, ➀ current work predominantly focuses on single-scale prediction and analysis, either coarse-grained phases or fixed temporal windows [36, 37]. Despite some improvements, this myopic perspective neglects the hierarchical and time-varying nature of surgical decision-making [38]. Additionally, attributed to the accumulation of inaccurate generations in foundation models like VLM and LLM and the increase in visual tokens caused by inputting past multiple frames, ➁ there is still significant room for improvement in the performance of existing architectures when dealing with surgical workflow analysis and prediction tasks. Additionally, the increasing input length from streaming input frames poses significant challenges for existing methods in terms of computation and memory, leading to performance degradation with the linear growth of temporal input [39]. Furthermore, the accumulation of errors in generative models hinders the effectiveness of existing methods in intelligent surgical problem solving [40, 41].
Several recent studies have incrementally addressed these limitations, though key challenges remain. GraSP [42] advances spatial scene understanding via pixel-level segmentation but overlooks temporal multi-scale dynamics. Similarly, SWAG [43] explores multi-step forecasting within single abstraction levels, yet its effectiveness is constrained by hierarchical dependencies. Meanwhile, alternative approaches, such as MAS, have been successfully leveraged across multiple dimensions of robotic workflows [44, 45], demonstrating the potential of multi-scale analysis for performance enhancement. However, existing efforts remain fragmented, and the trustworthiness of autonomous operations could be further improved through more advanced, integrated methodologies.
@@
In light of this, we propose IG-MC (Incremental Generation via Multi-agent Collaboration), a unified closed-loop framework for multi-scale temporal prediction. At its core, IG-MC introduces two key innovations: (1) an incremental generation mechanism that dynamically synthesizes predicted states and images across expanding temporal scales to maintain state-visual synchronization, thereby reducing error accumulation in long-horizon predictions and ensuring trustworthiness; and (2) a decision-driven multi-agent collaboration system comprising specialized VLM-based agents [46, 47] that hierarchically refine predictions from phases to steps while enforcing cross-scale consistency. Unlike existing methods that rely on historical frames, at each time step, the decision-making module (DM) of IG-MC predicts the state using only the current state-image pair, which then guides a visual generation module (VG) [48, 49, 50] to generate the next visual preview. Crucially, these innovations are embedded within a closed-loop design, where the state predictions of DM module and the visual outputs of VG module continuously inform and correct one another. This bidirectional interaction enables real-time error correction, interpretable forecasting, and high-fidelity coherence across both global and local scales.
To rigorously evaluate multi-scale forecasting in natural and surgical scene, we introduce the first Multi-scale Temporal Prediction (MSTP) benchmark, enhanced with synchronized multi-scale annotations across multiple temporal horizons and hierarchical state. Through extensive experiments, our framework demonstrates excellent performance across various metrics. After adding the plug-and-play modules (DM, VG), most metrics improve significantly.
@@ 2 Related Work
Surgical Workflow Recognition Automated recognition of surgical workflow phases and steps has been studied extensively. Early CNN-based multi-task methods [59] such as EndoNet learn phase recognition and instrument presence jointly from laparoscopic videos [60]. SV-RCNet further integrates a ResNet backbone with an LSTM for online workflow segmentation [61]. More recent temporal convolutional approaches [62], notably TeCNO, employ multi-stage dilated causal convolutions to refine surgical phase predictions iteratively and enforce smooth transitions [23]. Transformer-based models [63, 64] like MuST introduce multi-scale temporal encoding modules to capture short-, mid-, and long-term dependencies in surgical videos [65].
@@ 4.2 Evaluating the Efficacy of IG-MC (RQ1 & RQ4)
Obs 1. IG-MC maintains high performance as temporal scales increase. As shown in Fig 3, we can easily observe that the performance of introducing IG-MC remains at a high level over time. For example, when temporal scale is 5s, the Accuracy remains at 41.62% at time step 5. Compared with the effect of baseline 12% under the same conditions, it is significantly improved.
@@
Obs 3. IG-MC generalizes to diverse real-world scenarios, including general human action and surgical workflows. In general scenarios tested on the Action Genome (AG) dataset, IG-MC achieves an F1 score of 51.07 for joint attention-spatial-contacting state predictions at 10–20s scales, outperforming baselines by 10.24 points. In surgical contexts, the framework maintains 28.87% accuracy at 30s temporal scale with 5s incremental updates (24.95 $\uparrow$ compared to the baseline). This dual applicability underscores IG-MC’s potential as a unified framework for multi-scale prediction in both clinical and general embodied AI scenarios.
@@ 4.4 Performance in Various Scenarios (RQ3)
Obs 2. IG-MC achieves strong performance on hierarchical state scales, from coarse phases to fine steps On the phase-level state scale of MSTP-Surgery shown in Table 6, IG-MC achieves an F1 score of 56.11 when the temporal scale is 5s, compared to the baseline of 13.48. At the finer step-level scale, the framework improves recall by 31.78 points (from 7.94% to 39.72%) at the same temporal scale, showcasing its ability to model granular surgical actions while respecting hierarchical constraints. In cross-scale consistency tests, the joint phase-step prediction (P&S) achieves a Jaccard index of 20.06 with 5s temporal scale, much higher than the baseline of 1.57, confirming that IG-MC maintains coherence between abstract phases and concrete steps.
Obs 3. IG-MC adapts effectively to different incremental time scales, optimizing prediction granularity. When using a 1s incremental scale for 5s temporal predictions, IG-MC’s DM+VG setup achieves an accuracy of 42.40% on MSTP-Surgery, slightly below the DM-only baseline (50.80%) because of the quality of the pictures generated by VG. But with a 4.50-point recall improvement, it indicates the enhancement of step-level detail of the overall stability. At longer incremental scales (e.g., 5s for 30s temporal predictions), the framework retains 41.10% accuracy while improving recall by 20.18 points, demonstrating that coarser incremental updates can prioritize structural consistency without severe performance degradation. This flexibility allows IG-MC to trade off between computational efficiency and prediction granularity, making it suitable for real-time and planning-oriented scenarios alike.

### v3

@@ Abstract
Accurate temporal prediction is the bridge between comprehensive scene understanding and embodied artificial intelligence. However, predicting multiple fine-grained states of scene at multiple temporal scales is difficult for vision-language models. We formalize the Multi‐Scale Temporal Prediction (MSTP) task in general and surgical scene by decomposing multi‐scale into two orthogonal dimensions: the temporal scale, forecasting states of human and surgery at varying look‐ahead intervals, and the state scale, modeling a hierarchy of states in general and surgical scene. For instance in general scene, states of contacting relationship are finer-grained than states of spatial relationship. For instance in surgical scene, medium‐level steps are finer‐grained than high‐level phases yet remain constrained by their encompassing phase. To support this unified task, we introduce the first MSTP Benchmark, featuring synchronized annotations across multiple state scales and temporal scales. We further propose a novel method, Incremental Generation and Multi‐agent Collaboration (IG-MC), which integrates two key innovations. Firstly, we propose an plug-and-play incremental generation to keep high-quality temporal prediction that continuously synthesizes up-to-date visual previews at expanding temporal scales to inform multiple decision-making agents, ensuring decision content and generated visuals remain synchronized and preventing performance degradation as look‐ahead intervals lengthen. Secondly, we propose a decision‐driven multi‐agent collaboration framework for multiple states prediction, comprising generation, initiation, and multi‐state assessment agents that dynamically triggers and evaluates prediction cycles to balance global coherence and local fidelity. Extensive experiments on the MSTP Benchmark in general and surgical scene show that IG‐MC is a generalizable plug-and-play method for MSTP, demonstrating the effectiveness of incremental generation and the stability of decision‐driven multi‐agent collaboration. The code, model weights, benchmark can be found in https://github.com/jinlab-imvr/MSTP.
@@ 1 Introduction
In a parallel vein, recent advances in artificial intelligence [19, 20, 21, 22] have significantly improved analysis and prediction of healthcare robot, achieving promising performance in tasks such as phase recognition [23] and instrument-tissue interaction detection [24], through convolutional [25, 23], recurrent [26, 27, 28], and transformer-based [29, 30, 31] architectures, as well as video generation through diffusion models [32, 33]. Undoubtedly, these approaches made an initial breakthrough, enabling faithful and effective surgical problem-solving. However, ➀ current work predominantly focuses on single-scale prediction and analysis, either coarse-grained phases or fixed temporal windows [34, 35, 36, 37]. Despite some improvements, this myopic perspective neglects the hierarchical and time-varying nature of surgical decision-making [38]. Additionally, attributed to the accumulation of inaccurate generations in foundation models like VLM and LLM and the increase in visual tokens caused by inputting past multiple frames, ➁ there is still significant room for improvement in the performance of existing architectures when dealing with surgical workflow analysis and prediction tasks. Additionally, the increasing input length from streaming input frames poses significant challenges for existing methods in terms of computation and memory, leading to performance degradation with the linear growth of temporal input [39]. Furthermore, the accumulation of errors in generative models hinders the effectiveness of existing methods in intelligent surgical problem solving [40, 41].
Several recent studies have incrementally addressed these limitations, though key challenges remain. Across general scenes, long-video MLLMs and planners relax fixed windows and address streaming: Video-MME reveals persistent hour-scale reasoning gaps [42]; VideoStreaming and LongVLM bound token growth via memory propagation and hierarchical segment tokens [43, 44]; and HM-Diffuser targets extendable long-horizon planning with multiscale diffusion [45]. Nonetheless, many pipelines still rely on sparse keyframes or heavy compression, exhibit unfavorable memory/latency scaling, and accumulate rollout errors over long horizons. In surgical scenes, GraSP offers multi-granular supervision unifying pixel-level instruments/actions with step/phase labels [46], while SWAG anticipates minute-scale workflows with single-pass and autoregressive decoders [47]. Yet datasets remain procedure-specific and clip-centric, cross-scale dependencies are only partially modeled, and prospective clinical validation is limited. These gaps motivate an integrated, trustworthy multi-scale temporal predictor that couples short-horizon action accuracy with long-horizon task coherence under realistic streaming constraints.
@@
In light of this, we propose IG-MC (Incremental Generation via Multi-agent Collaboration), a unified closed-loop framework for multi-scale temporal prediction. At its core, IG-MC introduces two key innovations: (1) an incremental generation mechanism that dynamically synthesizes predicted states and images across expanding temporal scales to maintain state-visual synchronization, thereby reducing error accumulation in long-horizon predictions and ensuring trustworthiness; and (2) a decision-driven multi-agent collaboration system comprising specialized VLM-based agents [48, 49] that hierarchically refine predictions from phases to steps while enforcing cross-scale consistency. Departing from approaches that rely on frame histories, IG-MC uses a stateless DM that consumes only the current state–image pair to predict the next state, which then conditions the VG to synthesize the next visual preview [50, 51, 52]; removing temporal buffers reduces memory and speeds up training and inference. Crucially, these innovations are embedded within a closed-loop design, where the state predictions of DM module and the visual outputs of VG module continuously inform and correct one another. This bidirectional interaction enables real-time error correction, interpretable forecasting, and high-fidelity coherence across both global and local scales.
To rigorously evaluate multi-scale forecasting in general and surgical scene, we introduce the first Multi-scale Temporal Prediction (MSTP) benchmark, enhanced with synchronized multi-scale annotations across multiple temporal horizons and hierarchical state. Through extensive experiments, our framework demonstrates excellent performance across various metrics. After adding the plug-and-play modules (DM, VG), most metrics improve significantly.
@@ 2 Related Work
Workflow Recognition Automated recognition of surgical workflow phases and steps has been studied extensively. Early CNN-based multi-task methods [61] such as EndoNet learn phase recognition and instrument presence jointly from laparoscopic videos [25]. SV-RCNet further integrates a ResNet backbone with an LSTM for online workflow segmentation [62]. More recent temporal convolutional approaches [63], notably TeCNO, employ multi-stage dilated causal convolutions to refine phase predictions iteratively and enforce smooth transitions [23]. Building on these, Transformer-based models [64, 65] including Trans-SVNet[66], which uses a hybrid embedding aggregation strategy to fuse spatial and temporal cues, and MuST[37], which introduces multi-scale temporal encoding to capture short-, mid-, and long-term dependencies, further improve accuracy and temporal consistency on standard surgical video benchmarks.
Visual Generation Text to image diffusion achieves high fidelity and controllable synthesis through latent diffusion [67] and language conditioned models [68]. Spatial controls based on depth, edge, or segmentation provide precise guidance [69]. In surgical settings, endoscopic frame synthesis based on Stable Diffusion [70] and guided generators for segmentation and triplet constraints [71, 72] produce realistic frames that support augmentation and downstream analysis. Building on these image priors, video generation extends control over time. Open domain systems such as Sora improve long range temporal consistency [73]. Surgical variants including Endora [32], SurGen [74], and SurgSora [75] provide controllable previews aligned with surgical workflows for endoscopy and operative scenes.
@@ 4.2 Evaluating the Efficacy of IG-MC (RQ1 & RQ4)
Obs 1. IG-MC maintains high performance as temporal scales increase. As shown in Fig 3, we can easily observe that the performance of introducing IG-MC remains at a high level over time. For example, when temporal scale is 5s, the Accuracy remains at 42.06% at time step 5. Although it exhibits a decrease of 3.62% compared to time Step 1, its performance remains favorable.
@@
Obs 3. IG-MC generalizes to diverse real-world scenarios, including general human action and surgical workflows. In general scenarios tested on the Action Genome (AG) dataset, IG-MC achieves an F1 score of 51.07 for joint attention-spatial-contacting state predictions at 10–20s scales, outperforming baselines by 10.24%. In surgical contexts, the framework maintains 28.87% accuracy at 30s temporal scale with 5s incremental updates (24.95% $\uparrow$ compared to the baseline). This dual applicability underscores IG-MC’s potential as a unified framework for multi-scale prediction in both clinical and general embodied AI scenarios.
@@ 4.4 Performance in Various Scenarios (RQ3)
Obs 2. IG-MC achieves strong performance on hierarchical state scales, from coarse phases to fine steps On the phase-level state scale of MSTP-Surgery shown in Table 6 (See Appendix E), IG-MC achieves an F1 score of 56.11 when the temporal scale is 5s, compared to the baseline of 13.48. At the finer step-level scale, the framework improves recall by 31.78 points (from 7.94% to 39.72%) at the same temporal scale, showcasing its ability to model granular surgical actions while respecting hierarchical constraints. In cross-scale consistency tests, the joint phase-step prediction (P&S) achieves a Jaccard index of 20.06 with 5s temporal scale, much higher than the baseline of 1.57, confirming that IG-MC maintains coherence between abstract phases and concrete steps.
Obs 3. IG-MC adapts effectively to different incremental time scales, optimizing prediction granularity. When using a 1s incremental scale for 5s temporal predictions in Table 1, IG-MC’s DM+VG setup achieves an accuracy of 42.40% on MSTP-Surgery, slightly below the DM-only baseline (50.80%) because of the quality of the pictures generated by VG. But with a 4.50-point recall improvement, it indicates the enhancement of step-level detail of the overall stability. At longer incremental scales (e.g., 5s for 30s temporal predictions), the framework retains 41.10% accuracy while improving recall by 20.18 points, demonstrating that coarser incremental updates can prioritize structural consistency without severe performance degradation. This flexibility allows IG-MC to trade off between computational efficiency and prediction granularity, making it suitable for real-time and planning-oriented scenarios alike.
@@ 5 Conclusion
Acknowledgments and Disclosure of Funding
This work was supported by Ministry of Education Tier 2 grant, Singapore (T2EP20224-0028), and Ministry of Education Tier 1 grant, NUS, Singapore (23-0651-P0001).
@@ Appendix I Visual Prediction Metrics
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We introduce the MSTP Benchmark and the IG-MC framework with incremental generation and multi-agent collaboration, and highlight their effectiveness in enhancing multi-scale prediction accuracy and consistency compared to baselines.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: In this work, we systematically discuss the limitations of our research.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: This paper does not include experimental results related to theoretical aspects.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide the code necessary for replicating the studies described in this paper via an anonymous link, and we detail the experimental setup for the replication in the article itself.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: For the datasets disclosed in the article, we have provided information regarding their sources and origins.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: we have specified all the model details necessary to understand the results.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: In this paper, we have reported error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: In this paper, we provide detailed information about the experimental resources, including GPU configurations used in our studies.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The study presented in this paper conforms to the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We have provided the societal impacts of the work
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper does not address issues related to this aspect.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All creators and original owners of the assets used in our paper, such as code, data, and models, have been properly credited. We have explicitly mentioned the licenses and terms of use for each asset and have ensured full compliance with these terms throughout our research.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The research presented in this paper is not concerned with new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve experiments or research related to human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not address potential risks incurred by study participants.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: LLMs are only used for polishing the article and do not serve as important, original, or non-standard components of the core methods.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.